OpenStudy (anonymous):

Using a directrix of y = -2 and a focus of (2, 6), what quadratic function is created?

OpenStudy (anonymous):

f(x) = -1/8(x - 2)2 - 2 f(x) = 1/16(x - 2)2 + 2 f(x) = 1/8(x - 2)2 - 2 f(x) = - 1/16(x + 2)2 - 2

OpenStudy (anonymous):

we don't need to do much work since you have choices do you know what this looks like?

OpenStudy (anonymous):

ok good and not really /:

OpenStudy (anonymous):

i will graph the focus and directrix

OpenStudy (anonymous):



OpenStudy (anonymous):

the vertex is half way between the focus and the directrix, so it is at \((2,2)\)

OpenStudy (anonymous):



OpenStudy (anonymous):

ok i follow you

OpenStudy (anonymous):

then the equation is has to be either f(x) = 1/16(x - 2)2 + 2 or f(x) = 1/8(x - 2)2 - 2 because it opens up the distance between the vertex and the directrix is \(p=4\) so \(4p=16\) pick

OpenStudy (anonymous):

pick \[f(x)=\frac{1}{16}(x-2)^2+2\]

OpenStudy (anonymous):

here is the check http://www.wolframalpha.com/input/?i=parabola+y%3D1%2F16%28x+-+2%29^2+%2B+2

OpenStudy (anonymous):

ok thanks for all your help (:

OpenStudy (aum):

In f(x) = a(x-h)^2 + k, (h,k) is the vertex. Here the vertex is (2,2) and so the choice is f(x) = 1/16(x-2)^2 + 2

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ok thank you to @aum for the effort! (: